Minute of KAGRA commissioning meeting
- Date: 2018/8/7 11AM~12PM
- Attendance: Miyakawa, Aso, Yamamoto, Kokeyama
- eZuce: Kimura, Yokozawa, Uchiyama, Izumi, Haino, Darken
PSL, IOO (Kokeyama)
- PMC is influenced by KOACH filters.
- MZ has been locked with output RF modulations.
- RAM monitor can be measured at local tools. It will be connected to monitor system at the control room.
- FSS was recovered with new laser.
- IMMT2 had bad coil, it will be repaired.
- High power laser will be out around September in the PLS after safety. Beam dump will be installed 12/3 then real high power laser is provided to IFO.
Daily meeting
Night work
ALS
- 70-80% at optical table in the PSL room is done.
- Beam reached to POP table.
- One of laser is not healthy. Low power operation is temporarily OK. Spare will be purchased soon.
- Needs little more time for electronics.
- Trasmittance was 50-60% for optical fiber.
Buffle PD
- It was recovered all the PDs attached again to the black plate.
Others
KAGRA/Commissioning/Weekly/20180821/Minute (last edited 2018-08-21 12:04:07 by OsamuMiyakawa)